# Frustratingly Easy Zero-Day Audio DeepFake Detection via Retrieval Augmentation and Profile Matching

### v1

@@ Abstract
Modern audio deepfake detectors using foundation models and large training datasets have achieved promising detection performance. However, they struggle with zero-day attacks, where the audio samples are generated by novel synthesis methods that models have not seen from reigning training data. Conventional approaches against such attacks require fine-tuning the detectors, which can be problematic when prompt response is required. This study introduces a training-free framework for zero-day audio deepfake detection based on knowledge representations, retrieval augmentation, and voice profile matching. Based on the framework, we propose simple yet effective knowledge retrieval and ensemble methods that achieve performance comparable to fine-tuned models on DeepFake-Eval-2024, without any additional model-wise training. We also conduct ablation studies on retrieval pool size and voice profile attributes, validating their relevance to the system efficacy.  [^1] This study is supported by JST AIP Acceleration Research (JPMJCR24U3) and the New Energy and Industrial Technology Development Organization (NEDO, JPNP22007). This study was partially carried out using the TSUBAME4.0 supercomputer at the Institute of Science Tokyo.
Index Terms—  Audio DeepFake Detection, Retrieval Augmentation, Profile Matching.
1 Introduction
@@
With the advancement of deep neural networks (DNNs), modern audio deepfake detection (ADD) systems, or so-called spoofing countermeasures (CM), have achieved strong performance in both controlled and complex scenarios. Such progress is demonstrated through established evaluation frameworks like the evolving ASVspoof challenge series [1, 2, 3], as well as recent datasets including In-the-Wild [4], SpoofCeleb [5], and DeepFake-Eval-2024 (DE2024) [6]. Meanwhile, recent advances in self-supervised learning (SSL) audio feature extractors [7, 8] have significantly improved CM detection performance on these datasets, particularly when combined with backend classifiers such as AASIST [9, 10] and simple multi-layer perceptrons [11, 12].
Fig. 1: Demonstration of the vulnerability of deepfake detection models caused by the time lag between the zero-day attack period and model update. The time progression indicates natural time flow.
@@
However, current CMs primarily focus on empirical performance. Although they can be updated through fine-tuning, it demands significant computational resources and time, and thus often cause considerable delay in adapting to newly-emerged attacking algorithms, as illustrated in Figure 1. These algorithms, represented by not only research outputs, but also commercial services like ElevenLabs [^2] https://elevenlabs.io, may not be represented in currently-rolling CM training data during that period. Thus, they are called zero-day attacks and should ideally be addressed promptly without fine-tuning or additional training. This need for training-free ADD methodologies motivates the present study. We address the research question: Can we prevent zero-day spoofing attacks without any form of additional fine-tuning?
@@
We propose a simple yet comprehensive framework for handling unseen zero-day attacks using retrieval augmentation (RA) techniques. Our method maintains a knowledge base containing reference “seen” data, and addresses novel attacks through knowledge retrieval and ensemble methods based on k-nearest neighbors (KNN). The knowledge database stores both representation-level features and score-level predictions from a well-trained SSL-based CM. Additionally, we incorporate speaker and voice-related information to evaluate whether such profile matching can enhance the performance. Both the methods and their related models do not involve any additional re-training or fine-tuning, enabling prompt reaction and adaptation for robustness to the zero-day attacks.
While RA has been applied to ADD, existing methods significantly differ from our approach. [13] maintains a vector data pool and retrieves the vectors of the most closely-matched samples during inference to augment the detection process. [14] also uses KNN to select features for fine-tuning the backend classifier. However, apart from fine-tuning the SSL frontend for ADD, both work involve additional training of backend classifier for feature selection and decision making. Our proposed method employs RA by directly using existing massively-trained CMs, enabling more broadly applicable solution for advanced SSL-based CMs.
@@
Fig. 2: Outline of our methodology. The knowledge source data contains input recordings and ground truth binary labels. The lock symbols indicate that both the CM and profile feature extractors remain frozen without fine-tuning. Retrieved vectors can be either CM features, profile features, or both, depending on the strategy detailed in Section 2.3. Example number of retrieved utterances is $k=4$ in this outline.
2 Retrieval Augmentation Based on Knowledge Resources
2.1 Source data: DeepFake-Eval-2024
@@
We select DE2024 [6] as our source dataset for this study, as a representative of real-world deepfake threats with current generation techniques due to its sourcing methodology. DE2024 is a multi-modal dataset containing 56.5 hours of audio content collected from social media platforms and deepfake detection services in 2024. It includes over 40 languages from more than 80 web sources, with audio sampled at 44.66 kHz by default.
@@
In order to adapt and perform a fair comparison with the existing benchmark and up-to-date, widely used CMs, we implement the following adaptation schemes on DE2024: 1) Following the fine-tuning setup in [6], 60% and 40% of the original dataset is used for training and evaluation, respectively. Among the training data, we randomly picked 5% of them as a development set. The ratio of real to fake audios across all splits is approximately 39% vs. 61%, with minor variations observed across different durations; 2) To compare with the established baseline [^3] https://huggingface.co/nii-yamagishilab/wav2vec-large-anti-deepfake-nda, whose pre-trained SSL-based CM is acquired in this study, we segment both training and evaluation partitions into various lengths ranging from 4 seconds (as reported in [6]) to 50 seconds. The training partition is used to construct a knowledge base, maintaining the same real-to-fake data ratio. The evaluation partition serves as the query data, containing zero-day attacks. They are detailed in next subsection.
2.2 Knowledge Database Construction
@@
  • $s_{i}\in(0,1)$: This is the prediction score for knowledge ensemble. The scalar value is also obtained from the CM, where the SSL feature vector is passed through the linear layer of the CM, followed by a softmax to produce the score.
The detailed description of the SSL-based CM system and the profile feature extraction process is presented in Section 3.1.
2.3 Knowledge Retrieval & Ensemble
After constructing the knowledge database, we proceed to the retrieval and ensemble stage, shown in the bottom part of Fig. 2. Let us focus on the retrieval stage first. Given a query utterance $q$, we use the same SSL CM and profile extractor as constructing $\mathcal{K}$ to obtain the query CM feature vector $\mathbf{x}_{q,cm}$, query feature profile vector $\mathbf{x}_{q,prof}$, and prediction score $s_{q}$. Note that we do not have the binary ground truth label from the query side.
@@ 2.3 Knowledge Retrieval & Ensemble
We then fetch the neighbor labels, scores, and feature vectors from $\mathcal{K}$, denoting them as $Y_{q}=\{y_{i_{j}}\}_{j=1}^{k}$, $S_{q}=\{s_{i_{j}}\}_{j=1}^{k}$, and $X_{q}=\{\mathbf{x}_{i_{j}}\}_{j=1}^{k}$ respectively. With this retrieved set of data, we employ three simple and comprehensive ensemble strategies to aggregate the retrieved information into a final prediction:
@@
We try different combinations of retrieval and ensemble strategies as specified in Table 1, resulting in $3\times 3=9$ systems (S01 to S09).
3 Experimental Setup
3.1 Self-Supervised Learning Models
This section describes the two SSL-based models applied for both knowledge database construction and query inference.
@@ 3.1 Self-Supervised Learning Models
Countermeasure. We use an open-sourced SSL-based CM model [^4] https://huggingface.co/nii-yamagishilab/wav2vec-large-anti-deepfake-nda for this study. The model uses a Wav2Vec frontend [15] to extract feature vectors, followed by a backend classifier with adaptive pooling, a fully-connected MLP (without bias), and softmax operation for prediction scores. The model was post-trained [^5] As described in [11], post-training is an intermediate supervised training phase for SSL models, with specifically collected data and training schemes to adapt them for certain task. Here, the task is ADD. on approximately 74,000 hours of real and fake audio clips from various datasets, without any form of data augmentation. The model architecture and training are described in further detail in [11].
@@
Profile extractor. We employ vox_profile [^6] https://github.com/tiantiaf0627/vox-profile-release [16] to extract profile features using well-trained foundation models. We construct vectors by concatenating four attributes: 1) Age: a scalar representing age value in terms of 10 uniformed intervals from 0 to 100 years old777e.g., A value of 7 means an estimated age of 61-70 years old.; 2) Gender: a 2-dimensional one-hot vector for male/female classification; 3) Emotion: a scalar indicating the emotion trait (anger, fear, neutral, etc.) and 256-dimensional embedding vector extracted from the intermediate of the foundational model; 4) Voice quality: a 25-dimensional vector capturing five perceptual aspects: pitch, volume, clarity, rhythm, and voice texture, following the taxonomy defined in [17]. More details about how the representations are extracted can be found in [16].
3.2 Implementation & Evaluation
@@
Baselines. To demonstrate the effectiveness of RA against fine-tuning, we compare it with two baselines: 1) The post-trained model that returns $s_{q}$ from the DE2024 evaluation set; 2) A fine-tuned model from the post-trained one using the DE2024 training set (same data used for knowledge database). For fine-tuning, we use a learning rate of $1e^{-7}$ and other settings consistent with [11]. We fine-tune the post-trained model for 10 epochs and infer $s_{q}$ for evaluation, with notable improvements from the baseline, as reported in next section.
@@
Hyperparameters. The only hyperparameter is the number of utterances $k$ to be retrieved from the knowledge database. We range $k$ from 5 to 200 at several discrete values. For each system ID, we report results using the optimal $k$ value based on performance on the development set, and analyze the effect of $k$ through ablation study.
@@
Evaluation metric. We acquire two metrics in this study, equal error rate (EER), following standard practice in ADD, and accuracy with a fixed threshold of 0.5. We choose a fixed threshold instead of an optimized one from a development set of data because zero-day attacks involve new, unseen artifacts that make it difficult to derive proper thresholds from reigning data, similar to the challenges of model fine-tuning in practical scenarios.
4s 10s 13s 30s 50s
@@
System ID $\mathcal{N}_{q}$ $\hat{s}_{q}$ $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%)
S01 CM-only MV 200 9.24 78.63 200 7.94 81.25 200 8.09 82.54 20 7.69 82.78 20 9.04 86.27
S02 Ratio 20 18.40 81.98 20 15.37 84.63 20 15.82 84.46 100 15.54 83.50 20 13.21 85.51
S03 Avg 100 29.75 68.65 200 22.92 76.77 200 20.50 77.95 5 15.61 83.71 5 13.07 85.89
S04 Profile-only MV 100 12.30 78.26 50 12.33 79.15 20 12.68 80.33 20 14.09 75.03 10 12.80 80.34
S05 Ratio 5 19.92 79.14 5 17.43 78.37 5 18.04 80.33 5 17.21 79.52 10 12.80 80.34
S06 Avg 50 35.36 68.80 200 31.21 69.13 20 28.07 71.07 5 25.36 75.08 5 30.32 76.39
S07 Hybrid MV 200 7.62 82.15 100 6.72 85.51 100 7.48 85.56 200 7.16 84.07 5 8.21 83.82
S08 Ratio 10 14.98 82.63 10 12.49 85.54 10 12.44 86.03 10 11.42 87.13 10 11.68 87.49
S09 Avg 50 31.56 73.30 200 25.32 75.84 50 22.45 77.63 10 17.29 69.67 200 14.60 86.08
Baseline - - - 29.86 73.00 - 23.22 76.59 - 21.14 78.10 - 15.84 83.45 - 13.63 85.51
Fine-tuned - - - 14.10 86.97 - 12.79 88.52 - 11.65 88.71 - 12.03 86.67 - 12.10 83.07
Table 1: Performance of the proposed retrieval augmentation methods on DE2024 evaluation set. $k$ denotes the optimal number of retrieved utterances. Numbers in bold and underlined indicate the best and second-best results under each condition, respectively.
4 Results and Analysis
4.1 Main results
Results for different RA strategies are presented in Table 1. There are several key findings which are listed below.
Retrieval augmentation methods which directly incorporate class labels for knowledge ensemble can handle zero-day attacks as effectively as fine-tuned models. The proposed methods with majority voting (MV) and ratio-based ensemble consistently outperform the baseline and generally reach comparable performance to that of fine-tuned models. The lowest and second-lowest EERs are all achieved by the proposed methods using majority voting, respectively from S07 and S01888Here, it is worth noting that since there are only two possible values for MV, the EER computed is the average of false acceptance and miss rates.. While fine-tuned models demonstrate the highest accuracy on shorter durations, the proposed methods, especially those using ratio-based ensemble (S02, S05, S08), achieve comparable performance on longer durations. In particular, S08 achieves either second-highest or highest accuracy among all systems. However, score averaging for the ensemble often cannot outperform the baseline. The above findings demonstrate the efficacy of RA for ADD as a binary classification task. Future work may further this approach for multi-classification tasks such as spoof-aware speaker verification [18].
The profile information does not work well independently as retrieval method but is sufficient compensation for the CM features. Regarding the retrieval method, the analysis and results show that using CM-only or Hybrid for retrieval achieved comparable performance to that of the fine-tuned models. In contrast, using only profile feature vectors, while still outperforming the baseline, could not achieve comparable performance to the fine-tuned models. This may be because the foundation models used for profile feature extraction primarily store voice and speaker information rather than real/fake information, as mentioned in Section 3.1. However, the performance of Hybrid indicates that voice and speaker information can effectively complement CM features. In the next subsection, we investigate the individual usefulness of these components.
Apart from the above main findings, it may be also worth noting that audio duration has an impact on the relative efficacy of RA. The performance differences between our proposed methods, baseline, and fine-tuned models decrease with longer audio segments. This may be due to more informative embeddings, making models less sensitive to different strategies. Investigating these duration-dependent effects is a possible direction for future work.
4.2 Ablation study
In this section, we investigate two variables: the number of retrieved utterances $k$ from the knowledge database, and the effect of different attributes embedded in the profile feature vector.
The effect of number of retrieved utterances. We pick S07 and S08 for this ablation. We vary number of utterances retrieved from the knowledge $k$ from 5 to 200, as same in the main experiment. We present the results using one shorter (4 seconds) and one longer duration (30 seconds) cases. The performance of both systems measured by EER is presented in Fig. 3. For S07, the performance appears consistent. This might be due to the limited possible value of scores, since only binary values are used; meanwhile, S08 can only achieved best performance with $k=10$, and even though still comparable to the baseline, it has become worse. Such findings are invariant to the duration of the audio segments for different cases.

### v2

# Zero-Day Audio DeepFake Detection via Retrieval Augmentation and Profile Matching

@@ Abstract
Modern audio deepfake detectors built on foundation models and large training datasets achieve promising detection performance. However, they struggle with zero-day attacks, where the audio samples are generated by novel synthesis methods that models have not seen from reigning training data. Conventional approaches fine-tune the detector, which can be problematic when prompt response is needed. This paper proposes a training-free retrieval-augmented framework for zero-day audio deepfake detection that leverages knowledge representations and voice profile matching. Within this framework, we propose simple yet effective retrieval and ensemble methods that reach performance comparable to supervised baselines and their fine-tuned counterparts on the DeepFake-Eval-2024 benchmark, without any additional model training. We also conduct ablation on voice profile attributes, and demonstrate the cross-database generalizability of the framework with introducing simple and training-free fusion strategies.
I Introduction
@@
With the advancement of deep neural networks (DNNs), modern audio deepfake detection (ADD) systems, or so-called spoofing countermeasures (CM), have achieved strong performance in both controlled and complex scenarios. Such progress is demonstrated through established evaluation frameworks like the ASVspoof challenge series [12, 16], as well as recent datasets including SpoofCeleb [10] and DeepFake-Eval-2024 (DE2024) [2]. Meanwhile, recent advances in self-supervised learning (SSL) audio features [3, 8] have notably improved CM detection performance on these datasets, particularly when combined with backend classifiers such as AASIST [9, 15] and simple multi-layer perceptrons [7, 18].
@@
However, current CMs primarily focus on empirical performance. Although they can be updated through fine-tuning, it demands notable computational resources and time, and thus often cause considerable delay in adapting to newly-emerged attacking algorithms, as illustrated in Figure 1. These algorithms are represented by not only research outputs, but also commercial services like ElevenLabs [^1] https://elevenlabs.io. They may not be represented in currently-rolling CM training data during that period. Thus, they are called zero-day attacks and should ideally be addressed promptly without fine-tuning or additional training. This need for training-free ADD methodologies motivates the present study. We address the research question: Can we defend zero-day audio Deepfake attacks promptly without any form of additional fine-tuning?
@@
To answer this question, we propose a simple and comprehensive framework against zero-day attacks, utilizing retrieval augmentation (RA). It utilizes a knowledge database containing reference seen data, and addresses the attacks through knowledge retrieval and ensembling based on k-nearest neighbors (k-NN). The knowledge database stores both representation-level features and score-level predictions from a well-trained SSL-based CM. Additionally, we incorporate speaker and voice-related information extracted from specialized foundation models, and seek whether leveraging such profile matching method can further enhance the detection performance. Both the methods and their related models do not involve any additional re-training or fine-tuning by themselves, enabling prompt reaction and adaptation for robustness to the zero-day attacks. Furthermore, in order to validate and enhance the generalizability of the framework, we introduce AI4T [4] as an alternative evaluation dataset, and extend the framework with two simple, training-free fusion strategies.
Figure 1: Demonstration of the vulnerability of deepfake detection models caused by the time lag between the zero-day attack period and model update. The time progression indicates natural time flow.
@@
Figure 2: Outline of the methodology. The knowledge source data contains input recordings and ground truth binary labels. The lock symbols indicate that both the CM and profile feature extractors remain frozen without fine-tuning. Retrieved vectors can be either CM features, profile features, or both, depending on the strategy detailed in Section II-C. Example number of retrieved utterances is $k=4$ in this outline.
While RA has been applied to ADD, existing methods substantially differ from the approach proposed in this study. [11] maintains a vector data pool and retrieves the vectors of the most closely-matched samples during inference to augment the detection process. [13] employs k-NN to select features for fine-tuning the backend classifier. However, apart from fine-tuning the SSL frontend for ADD, both works involve additional training of backend classifier for feature selection and decision making. This study employs RA by directly using existing CMs that has been trained on massive amount of ADD-related datasets, enabling more broadly applicable solution for advanced SSL-based CMs. For completeness, we also repeat the same proposed method via knowledge database with a model fine-tuned on DE2024 training data, providing supplementary insights and contextual positioning of our approach.
II Knowledge-Based Retrieval Augmentation
II-A Source data: DeepFake-Eval-2024
@@
DE2024 [2] serves as the source dataset for this study, as a representative of real-world deepfakes. It contains 56.5 hours of audio content collected from various social media platforms and deepfake detection services, and includes over 40 languages from more than 80 web sources.
@@
In order to adapt and perform a fair comparison with the existing, widely used CMs, we implement the following adaptation schemes on DE2024: 1) Following the fine-tuning setup in [2], 60% and 40% of the original dataset is used for training and evaluation, respectively. Among the training data, we randomly picked 5% of them as a development set. The ratio of real to fake audios across all splits is approximately 39% vs. 61%, with minor variations observed across different durations; 2) To compare with the established baseline, whose pre-trained SSL-based CM is acquired in this study, we segment both training and evaluation partitions into various lengths ranging from 4 seconds (as reported in [2]) to 50 seconds. The training partition is used to construct the knowledge database, maintaining the same real-to-fake data ratio. The evaluation partition serves as the query data, containing zero-day attacks. They are detailed in next subsection.
II-B Knowledge database construction
@@
  • $s_{i}\in(0,1)$: This is the prediction score for knowledge ensemble. The scalar is also obtained from the CM, where the SSL feature vector is passed through the linear layer of the CM, followed by a softmax to produce the score.
The detailed description of the SSL-based CM and the profile feature extraction process is presented in Section III-A.
II-C Retrieval augmentation
After constructing the knowledge database, we proceed to performing RA, shown in the bottom part of Fig. 2. Firstly, at the retrieval stage, given a query utterance $q$, we use the same SSL-based CM and profile extractor as constructing $\mathcal{K}$ to obtain the query CM feature vector $\mathbf{x}_{q,cm}$, query profile feature vector $\mathbf{x}_{q,prof}$, and prediction score $s_{q}$. Note that we do not have the binary ground truth label from the query side.
@@ II-C Retrieval augmentation
We then proceed to the ensemble stage, fetching the neighbor labels, scores, and feature vectors from $\mathcal{K}$, denoting them as $Y_{q}=\{y_{i_{j}}\}_{j=1}^{k}$, $S_{q}=\{s_{i_{j}}\}_{j=1}^{k}$, and $X_{q}=\{\mathbf{x}_{i_{j}}\}_{j=1}^{k}$ respectively. We employ three simple ensemble strategies to aggregate the retrieved information into a final prediction:
@@
The ensembled score is the output of RA. We try different combinations of retrieval and ensemble strategies as specified in Table I, resulting in $3\times 3=9$ systems (S01 to S09).
4 seconds 10 seconds 13 seconds 30 seconds 50 seconds
[…]
System $\mathcal{N}_{q}$ $\hat{s}_{q}$ $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%)
S01 CM-only MV 20 15.47 86.46 20 13.96 87.81 10 13.15 88.65 20 13.76 88.00 200 16.96 85.98
S02 Ratio 20 14.62 86.43 100 14.09 86.70 20 12.98 87.96 20 11.21 88.92 20 11.59 87.49
S03 Avg 5 27.27 75.73 5 21.91 78.90 200 20.25 79.16 200 16.56 81.21 100 15.35 82.41
S04 Profile-only MV 100 26.07 78.26 50 24.84 79.15 20 23.02 80.33 20 23.86 79.26 10 23.40 80.34
S05 Ratio 100 25.35 78.22 5 23.59 78.37 5 22.18 79.16 5 22.17 79.52 10 22.75 78.55
S06 Avg 50 35.36 68.80 200 31.21 69.13 20 28.06 71.07 5 25.39 75.08 5 30.42 76.39
S07 Hybrid MV 200 15.70 86.38 100 14.69 88.01 100 14.63 87.96 20 13.08 89.07 20 13.87 88.52
S08 Ratio 20 14.78 86.84 10 12.81 87.73 10 12.81 87.52 20 11.25 89.38 20 10.92 88.90
S09 Avg 50 33.19 69.20 200 31.37 69.35 50 27.93 71.89 100 31.19 70.17 200 32.77 69.61
B00 - - - 27.76 75.15 - 22.21 78.95 - 20.36 79.14 - 16.67 81.00 - 14.92 82.22
B01 - - - 12.52 87.69 - 11.11 90.95 - 10.21 90.68 - 9.53 90.91 - 9.96 90.53
TABLE I: Performance of the proposed RA methods on DE2024 evaluation data. $k$ denotes the optimal number of retrieved utterances. Numbers in bold and underline indicate the best and second-best results under each condition, respectively.
III Experimental Setup
III-A Self-supervised learning models
@@
Countermeasures. We use an open-sourced SSL-based CM model [^2] https://huggingface.co/nii-yamagishilab/xls-r-2b-anti-deepfake for this study. The model uses a Wav2Vec2.0 frontend [1] to extract feature vectors, followed by a backend classifier with adaptive pooling, a fully-connected MLP (without bias), and softmax operation for prediction scores. The model was post-trained [^3] As described in [7], post-training is an intermediate supervised training phase for SSL models, with specifically collected data and training schemes to adapt them for certain task. Here, the task is ADD. on approximately 74,000 hours of real and fake audio clips from various datasets. Further details of model architecture and training are described in [7].
@@
Profile extractor. We employ vox_profile [^4] https://github.com/tiantiaf0627/vox-profile-release [6] to extract profile feature vectors using well-trained foundation models. We construct the features by concatenating four attributes: 1) Age: a scalar representing age value in terms of 10 uniformed intervals from 0 to 100 years old555e.g., A value of 7 means an estimated age of 61-70 years old.; 2) Gender: a 2-dimensional one-hot vector for male/female classification; 3) Emotion: a scalar indicating the emotion trait (anger, fear, neutral, etc.) and 256-dimensional embedding vector extracted from the intermediate of the foundational model; 4) Voice quality: a 25-dimensional vector capturing five perceptual aspects: pitch, volume, clarity, rhythm, and voice texture, following the taxonomy defined in [5]. More details about how the representations are extracted can be found in [6].
III-B Implementation & evaluation
@@
Baselines. To demonstrate the effectiveness of RA against fine-tuning, we compare it with two baselines: 1) The post-trained model that returns $s_{q}$ from the DE2024 evaluation data; 2) A fine-tuned model from the post-trained one using the DE2024 training data (same data used for knowledge database of RA). For fine-tuning, we use a learning rate of $1e^{-7}$ and other settings consistent with [7]. We fine-tune the post-trained model for 10 epochs and infer $s_{q}$ for evaluation, with notable improvements from the baseline, as reported in next section. The baseline and fine-tuned models are denoted as B00 and B01, respectively in the rest of this paper.
@@
Hyperparameters. The only hyperparameter is the number of utterances $k$ to be retrieved from the knowledge database. We range $k$ from 5 to 200 at several discrete values. For each system, we report results using the optimal $k$ value based on performance on the development set.
@@
Evaluation data. For validating the proposed RA framework itself, we use the DE2024 evaluation set under the corresponding various duration conditions. For the extension covered in Section V, we additionally acquire AI4T [4], a social-media audio Deepfake corpus with 13 hours of speech in 8 languages. We follow the official protocol described in [4] [^6] https://github.com/davidcombei/AI4T/tree/main to construct the evaluation data. This dataset differs substantially from DE2024 in its attacking methods, which makes it suitable for benchmarking generalizability.
Evaluation metrics. We acquire two metrics: 1) Equal error rate (EER), following standard practice in ADD; 2) Accuracy with a fixed threshold of 0.5. A fixed threshold is used instead of one optimized on a development set because zero-day attacks may involve new, unseen artifacts, making it difficult to derive reliable thresholds from existing data, similar to the limitations of model fine-tuning in real-world scenarios.
IV Results and Analysis
IV-A Main Results
Results for different RA strategies are presented in Table I. There are several key findings which are listed below.
Retrieval augmentation methods which directly incorporate class labels for knowledge ensemble are effective to zero-day attacks, being comparable to fine-tuned models. The proposed methods with MV and ratio-based scoring ensemble consistently outperform B00 in terms of accuracy. Among them, the ratio-based ensemble with a hybrid retrieval strategy (S08) achieves performance comparable to B01, having either the 2nd best EER or accuracy (or both) across all duration conditions. This indicates the effectiveness of the proposed information acquisition mechanism. In terms of EER, while the best results are generally achieved by the fine-tuned model, the ratio-based fusion using either the CM feature alone or both features (S02, S08) achieves nearly comparable performance, with the gap further narrowing for longer durations. Meanwhile, score averaging often fails to surpass the baseline. The above findings demonstrate the efficacy of RA for ADD as a binary classification task.
The profile information does not work well independently, but is sufficient compensation for the CM features. Regarding the retrieval method, the analysis and results show that using CM-only or Hybrid for retrieval, with Ratio as ensemble method (S02, S08), achieved comparable performance to that of the fine-tuned models. S08 reaches relatively more stably comparable performance against B01 across all duration conditions, which incorporates the profile information. In contrast, using only profile features (S04-S06) do not return as promising performance. This is likely because the foundation models used for profile feature extraction mainly capture voice and speaker characteristics rather than real/fake cues, as discussed in Section III-A. Nevertheless, the strong performance of S08 suggests that speaker- and voice-related information can effectively complement CM-based representations. Section IV-B examines the individual contribution of these components.
Apart from the above main findings, it may be also worth noting that audio duration may have an impact on the relative efficacy of RA. The relative performance gap between (especially) B00, the best of the proposed methods, and B01 decrease with longer audio segments. This may be due to more informative embeddings, making models less sensitive to different strategies. Investigating these duration-dependent effects is a possible direction for future work.
4 seconds 30 seconds
@@ IV-A Main Results
S08 14.78 86.84 11.25 89.38
w/o Age & Gender 14.53 87.06 12.06 88.76
w/o Emotion 14.48 86.88 12.15 88.76
w/o Voice Quality 16.14 85.54 13.07 87.05
TABLE II: Probing attributes from profile feature extractor (w/o: without). Number of retrieved utterances here is $k=20$.
IV-B Ablation on voice profile attributes
Earlier results shown in last section demonstrates that the profile features are good compensation. It is then natural to additionally analyze, among the several attributes involved in constructing the profile features, which ones drive to positive results more than others. We use S08 to compare the attributes by probing them in the profile features: age & gender, emotion, and voice quality. We remove each attribute individually to assess its contribution to performance.